OpenStudy (anonymous):

If f(x) = ∣(x2 − 6)(x2 + 2)∣, how many numbers in the interval 1 ≤ x ≤ 2 satisfy the conclusion of the mean value theorem?

OpenStudy (freckles):

\[f'(c)=\frac{f(2)-f(1)}{2-1}\] is what you need to solve

OpenStudy (anonymous):

I know that but I understand that you need the calculate the derivative of the equation and that is what i'm confused about and also how to solve to find the values of c.

OpenStudy (freckles):

\[f(x)=|(x^2+2)(x^2-6)| \\ x^2+2 \ge 0 \text{ for all real x } \\ f(x)=(x^2+2)|x^2-6| \\ \text{ we only care what happens with f when x is between 1 and 2 } \\ x^2-6=0 \text{ when } x=\pm \sqrt{6}\] so our numbers occur between the two zeros (or are interval is a subinterval of (-sqrt(6),sqrt(6)) so we need to find out what is going on between there

OpenStudy (freckles):

like is the function positive or negative on that interval

OpenStudy (freckles):

  i flipe the y values that corresponded to the x-values between -sqrt(6) and sqrt(6)